OpenStudy (anonymous):

I have a problem that I have to solve using square root property 5x^2-3=0

OpenStudy (anonymous):

5x^2=3 x^2=3/5 x=+ or - sqrt(3/5)

OpenStudy (anonymous):

it needs to be a solution set

OpenStudy (anonymous):

{-sqrt(3/5),sqrt(3/5)} if that's what you mean.